OpenStudy (anonymous):

square root bye division method of: 1)2304

OpenStudy (anonymous):

help plzzzzzz

OpenStudy (raden):

2+3+0+4 = 9 so, it means that number divisible by 9 look at the last two digit ... 04 it means that number divisible by (2^2 = 4) because divisible by 9 and 4, therefore that number also divisible by (9 * 4 = 36) just check it : 2304/36 = 64 in other words 2304 = 64 * 36 so, sqrt(2304) = sqrt(64 * 36) = sqrt(64) * sqrt(36) = 8 * 6 = 48

OpenStudy (anonymous):

square root by division arcane procedure is outlined here http://www.freeganita.com/en/nt/1_5.htm

OpenStudy (anonymous):

ok thnks

OpenStudy (anonymous):

:)